If the kinetic energy of an electron in the ground state hydrogen atom is 2.2 x 10-18 J, calculate its momentum


1
Expert's answer
2020-07-01T15:42:02-0400

The relationship between the momentum and kinetic energy of an electron is

"p=\sqrt{(2mK)}=\sqrt{(2(9.11*10^{-31})(2.2*10^{-18 }))}"

"=2.00*10^{-24} kg.m\/s"